HUTP-91/A000

This is the Template

Author One and Author Two^{†}

Lyman Laboratory of Physics Harvard University Cambridge, MA 02138

This sample is produced by the file ‘sample.tex’ in the directory tex$macros and is intended to exhibit the conventions used in the file ‘harvmac.tex’ also in this directory, as inputted above. (You do not need to copy tex$macros:harvmac.tex into your own directory, nor do you need to explicitly specify its directory as long as there is no harvmac.tex in your default directory.) The vskips, etc., on this title page have been chosen to give a reasonable looking HUTP style preprint in either full-sized or reduced format (selected by typing b or ℓ in response to the big/little query from the macro). Note that the ‘Title’ and ‘Date’ macros should be used together to choose fonts and horizontal size appropriately, suppress first page number, etc., for ‘big’ and ‘little’ formats with title page. One may most easily start a new paper by making a copy of this sample and then replacing the dummy text with something potentially more interesting. It is useful to read pp. 1–54, 127–201 of Knuth’s Texbook.

4/91

† (author@huhepl.hepnet, @huhepl.bitnet, or @huhepl.harvard.edu)

1. Introduction

Sections are begun using the macro \newsec as above. For papers without separate sections, leave out \newsec and equations will automatically be numbered consecutively (without section numbers).

References are done using the macro \ref as follows[1][2]. The labels on the references[1] and [2] (which include the brackets, note the second backslash to provide a space) can then be used to refer back to them[1]. (note no extra backslash needed before period[2], or before comma.) Here is how to put two papers in one reference[3]. Note the use of the macro \semi to separate them (the macro \nref can be used to suppress the immediate appearance of the bracketed reference number in the text if desired[4]–[6]). The references are listed in the file refs.tmp, and can be of arbitrary length.

The macro \foot is used for footnotes.^{1} They also get automatically numbered^{2} throughout the paper.

To do equations we use the macro \eqn

R^{a}_{b} = dω^{a}_{b} + ω^{a}_{c} ∧ ω^{c}_{b}  (1.1)

and

ω^{a}_{b,μ} ≡ e^{a}_{ν} ∇_{μ} e_{b}^{ν} = −e_{b}^{ν} ∇_{μ} e^{a}_{ν}.  (1.2)

The equations are then referred to (see (1.1) or (1.2)) using the label (which includes the parentheses). In draftmode, the equation labels appear in the right margin next to the equation.

2. Body

Ordinarily one uses equations with internal eqaligns:

\begin{matrix}d = 2 + 2 = 4 \\ d = 2 × 2 = 4. \end{matrix}  (2.1)

(How could Nature fail to avail herself of such a fine opportunity to base herself on the remarkable numerical coincidence embodied in eqn. (2.1)?) To specify equation numbers in aligned equations and/or equations specified with a’s, b’s, etc., just tag the equation numbers first using the macro \eqnn or \eqna and

then use \eqalignno

(∫_{−∞}^{∞} e^{−x^{2}} dx)^{2} = something  (2.2a)

= something else

= π  (2.2b)
= 2π/2.  (2.3)

We can then easily refer back to (2.3) or (2.2a) or (2.2a, b) or even (2.2a–c) (\eqna defines a label which takes an argument {a, b, …} when used to refer back; it could be used instead of \eqnn if one remembers to type {} after the label when referring back, e.g., (2.2) ). (Note also that \nref and \eqnn can always be used to simulate macro packages that require a separate macro to define labels in advance of using them.)

To list the references, use the macro \listrefs as at the end of this file^{3}.

3. Conclusion

To call Tex on a source file filename.tex, issue the command

$ tex filename

which creates a file filename.dvi (and also a filename.log file). The macro package will ask you whether you want “big” output, ordinary typewritten pages, or “little,” reduced two-column pages suitable for preprinting. (Default is for postscript laserprinters. In order to print out in reduced format on qms printer at HEPL, you need to specify \qms after inputting the macros.)

2 3.1. At Lyman

1 like this 2 consecutively

3 In draftmode reference labels appear in left margin next to reference.

After eliminating any tex errors in your manuscript, issue the command

$ lw filename

to convert the .dvi file to a postscript file that is then automatically sent to the lyman laserwriter. (To process only selected portions of a .dvi file, type instead $ dvips filename — then type i for interactive [or ? for help at any time]. $ lw filename.ps with explicit .ps extension then sends the .ps file to the Lyman laserwriter.)

3.2. At HEPL

(at hepl, use instead $ qmstex filename

to convert the .dvi file to a .qms file that is then automatically sent to the hepl laserprinter). Before logging off, it is probably preferable to get rid of any .ps, .dvi, .qms, and .log files.

Acknowledgements

This work was supported in part by various colleagues. We also thank NSF contract PHY-82-15249 for discussions.

Appendix A.

Everything same as before, now equations appear as (A.1), etc. Appendix A has no title.

Appendix B. Enhancements added, 11/88

For fancier treatment of multiple references, say [4–6], [2,4], or even [1,2,3– 6]. (The arguments have to be previously defined references, so use \nref before for any first occurrences.) To strip off the brackets from a single reference for some reason, use \xref to give 1 instead of [1].

To refer to equations (or references) that occur much later in the paper, add \writedefs after inputting macros. This redefines the equation and reference macros so they write all the label definitions to a file labeldefs.tmp. Then tex a second time with a line inputting this file. To start from the middle of a paper, tex the first part with \writedefs and \writestop (the latter adds to labeldefs a line with the pagenumber, section number, etc., at the point \listrefs was called). Inputting labeldefs.tmp then starts the continuation where the previous left off, and allows referring to any of the previously defined reference and equation labels.

The subsection option in section 3 above is also new.

Another macro, \footend, redefines \foot so that it writes the footnotes in a file foots.tmp, which is then automatically called in by \listrefs before the references. For journals that prefer the footnotes at the end, one can thus add \footend after inputting harvmac, with no other changes necessary.

A new macro \fig (also \nfig, \figs, \xfig) works for figures just as does \ref. The label is defined to include the abbreviation fig. For papers with thousands of figures, the macro fig. 1. This works exactly as does [7], with the label defined to include the abbreviation ‘fig. ’. The macro \nfig suppresses immediate appearance of ‘fig. x’, so that one can instead say, for example, figs. 2,3 using the macro \figs. \xfig strips off the ‘fig. ’ to leave just the figure number, i.e. 1, if necessary for some reason. The figure captions are written in a file figs.tmp that is called in by saying \listfigs after \listrefs. (Note that the formerly used macro \figures need not be used if \fig is instead used in conjunction with \listfigs.)

4 Appendix C. Enhancements added, 6/89

A macro \seclab sets a label equal to the current section number, useful for referring back to the section, e.g. as used in section 3 above. Similarly, \subseclab sets a label equal to the current section.subsection, as in the definition of subsection 3.2 in the “at HEPL” subsection above. (Using \writedefs, these definitions will appear in labeldefs.tmp so can be used as well to refer forward. In draftmode, the definitions appear in the right margin.)

A macro \writetoc writes a file toc.tmp that contains section and subsection numbers and headings with the pagenumbers on which they appear. The macro \listtoc writes the heading Contents and inputs the file toc.tex. It is evidently necessary to tex in two passes, renaming toc.tmp to toc.tex after the first.

Appendix D. Enhancements added, 1/91

A macro \lref (same syntax as \ref and \nref) defines the reflabel to be a macro that calls \ref on itself. This allows making a list of references via \lref (in any order) at the beginning of a file and only those used in the text will actually appear (automatically listed in correct order of first appearance by \listrefs). (In principle \lref can be used to compile a long standard reference file to input, although this might eventually cause problems due to clogging up tex memory with too many unused definitions.)

\refs is defined to not display reference labels that appear in succession (i.e. without intervening delimiter), so that e.g. [8–10] can expand the middle reference [9] if it hasn’t already appeared. \refs will also accept reflabels that are not previously defined so that it is possible to tex preliminary versions without inputting full reference list (For this mode, even individual refs, e.g. [11], need to be called via \refs.)

The macro \sequentialequations inserted at beginning causes equations to be numbered sequentially (1)–(n) throughout paper (rather than with section numbers (2.1)…).

References

[1] John Q. Physicist, Phys. Rev. D20 (1980) 1234.

[2] Author One and Author Two, Nucl. Phys. B52 (1975) 439.

[3] L. Alvarez-Gaumé and P. Ginsparg, “The Structure of Gauge and Gravitational Anomalies,” Harvard preprint HUTP-84/A016, to appear in Ann. Phys.; L. Alvarez-Gaumé and P. Ginsparg, Nucl. Phys. B243 (1984) 449.

[4] Seminal Paper.

[5] Moderately Interesting Follow-ups.

[6] Completely Useless Ripoffs.

[7] For a compendium of references missing here, see M. Green, J. Schwarz, and E. Witten, Superstring theory, Vols. I,II, Cambridge Univ. Press (1987).

[8] extra ref one.

[9] extra ref two.

[10] extra ref three.

[11] Another reference.

6

Figure Captions

Fig. 1. This is the figure caption Fig. 2. second dummy figure Fig. 3. third dummy figure 8